OpenStudy (notsosmart614):

which of the following is the probability of drawing a black card or a face card from a standard deck of cards

OpenStudy (notsosmart614):

1/2, 19/26, 3/26, 8/13

OpenStudy (benthicflare):

Standard deck of cards has 52 cards. Half of the cards are red (26), half of them are black (26) In the whole deck there are 4 suits (spades, hearts, diamonds, clubs). In each suit there are 3 face cards (Jack, Queen, King). There are therefore 12 face cards in total in the deck (half of which are red face cards, the other half are black face cards). Are you able to calculate from here?

OpenStudy (notsosmart614):

Can you help a little more im a little confused

OpenStudy (benthicflare):

Okay so the probability of drawing a black card would be 26/52 = 1/2 The probability of drawing a face card would be 12/52 = 3/13 Are you able to calculate the probability of drawing a black card OR a face card? What we know from here is that the probability definitely has to be more than 1/2 or 3/13.